OpenStudy (anonymous):

solve by the elimination method , 5x+6y=4 10x+12y=8 i got that there is no solution am i right?

OpenStudy (anonymous):

there are infinite solutions

OpenStudy (anonymous):

because if you multiply the first equation straight across by 2, you get the second equation exactly. so they are the same line

OpenStudy (anonymous):

so am i right ?

OpenStudy (anonymous):

sattelite can you help me really fast check some problems please

OpenStudy (anonymous):

you are not right

OpenStudy (anonymous):

sure i can check for you. what problems?

OpenStudy (anonymous):

just looked at the last one you posted. i wrote a solution

OpenStudy (anonymous):

difference between 'no solution ' and 'infinite solution' if second equation was 10x+12y=9 then no soloution

OpenStudy (anonymous):

solve the system of equations by graphing then classify the system as consitant or inconsistant and as dependent or independent 5x-5y=-15 5y-5x=15

OpenStudy (anonymous):

so its infinity solutions is that the answer